OpenStudy (niahloves):

Help! medal and fan Which statements are true according to the order of operations? Choose exactly two answers that are correct. A. 32 + 8 ÷ 4 = 40 ÷ 4 = 10 B. 4 + 12 ÷ 2 = 4 + 6 = 10 C. (20 – 5) • 3 = 15 • 3 = 45 D. 12 + (6 ÷ 3) = 18 ÷ 3 = 6

OpenStudy (anonymous):

b

OpenStudy (niahloves):

it's says two answers

OpenStudy (niahloves):

@texaschic101

OpenStudy (anonymous):

c

OpenStudy (anonymous):

b and c or b and a

OpenStudy (texaschic101):

b and c are true

OpenStudy (texaschic101):

because A is : 32 + 8 / 4 = 32 + 2 = 34 and D is : 12 + 6/3 = 12 + 2 = 14